# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ Appendix F More Ablation Study Results
F.1 Separate Tuning V.S. Unified Tuning
Here, we present the performance of separate tuning and unified tuning across the remaining 9 tasks.
Figure 10: Additional results from the ablation study comparing our unified instruction tuning to separate training.
As show in Figure 10, We observe that tasks like solvent prediction, catalyst prediction, and experiment procedure prediction continue to gain significant improvements from unified training. However, we also notice that some tasks perform worse after unified training compared to individual training, particularly those related to regression tasks, such as yield regression, logP prediction, and complexity prediction. While we see some modest gains in SCF prediction and weight prediction, there is a noticeable performance degradation in other tasks, especially in the complexity prediction task, where unified training causes a substantial drop in performance.
Regression tasks differ significantly from SELFIES generation and text description tasks in that the model must, after understanding the molecule, directly and accurately predict the specific numerical value of a property. Even though existing literature suggests that LLMs are capable of performing regression, accurately predicting the next number remains a fundamentally different challenge from current tasks, and potentially introduces significant conflicts.
Perhaps incorporating additional information, such as the process of deriving a specific number (e.g., a chain of thought), or using an additional regression head, could help alleviate these conflicts before making an accurate numerical prediction.
Appendix G Further Details on Efficiency Analysis
Figure 11: A comprehensive efficiency evaluation of our Omni-Mol, built on LLaMA 3.2-1B, compared to the state-of-the-art baseline. We conduct all experiments three times and compute the average result.
To ensure a fair comparison of computational efficiency with existing methods, we evaluate Omni-Mol across four key metrics: inference time, memory consumption, token generation speed (tokens per second), and GFLOPs. All comparative experiments are conducted on a single standalone NVIDIA A100-80G GPU to ensure consistency in inference evaluation. As illustrated in Figure 11, Omni-Mol achieves significantly lower inference latency, reducing it by up to 65% compared to PRESTO. This substantial efficiency gain enables faster response times, making Omni-Mol particularly well-suited for high-throughput molecular modeling tasks. Moreover, Omni-Mol demonstrates optimized memory utilization, consuming 28% less memory, which enhances its feasibility for deployment on resource-constrained hardware without compromising performance. Additionally, Omni-Mol achieves a 1.78$\times$ higher token generation rate, ensuring faster sequence generation and significantly improving usability in real-world applications that require rapid molecular property predictions and synthesis planning. Finally, Omni-Mol effectively reduces computational cost by 41% in terms of GFLOPs, striking a favorable balance between model complexity and inference speed. These advantages make Omni-Mol a scalable, cost-efficient, and high-performance solution for large-scale molecular modeling and chemistry-related AI applications.
@@ Appendix G Further Details on Efficiency Analysis
Appendix H Further Details on Theoretical Analysis
H.1 Proof of Theorem 3.1
Proof of Theorem 3.1.
For $F_{\text{general}}^{(n)}$ and $F_{\text{general}}^{(m)}$ with $m>n$, we have
$F_{\text{general}}^{(m)}$ $=F_{1}\cap F_{2}\cap\cdots\cap F_{n}\cap F_{n+1}\cap\cdots\cap F_{m}$ (18)
$=F_{\text{general}}^{(n)}\cap F_{n+1}\cap\cdots\cap F_{m}$
Since intersection of sets can only become smaller or remain the same as more sets are intersected, therefore
$F_{\text{general}}^{(m)}\subseteq F_{\text{general}}^{(n)}$ (19)
If $\forall i,j$, $n<i,j<m$ and $F_{i}\neq F_{j}$ when $i\neq j$, then
$F_{\text{general}}^{(n+1)}=F_{\text{general}}^{(n)}\cap F_{n+1}\subsetneq F_{\text{general}}^{(n)}$ (20)
$F_{\text{general}}^{(n+2)}=F_{\text{general}}^{(n+1)}\cap F_{n+2}\subsetneq F_{\text{general}}^{(n+1)}\subsetneq F_{\text{general}}^{(n)}$
$\cdots$
$F_{\text{general}}^{(m)}=F_{\text{general}}^{(n+(m-n-1))}\cap F_{n+(m-n)}\subsetneq F_{\text{general}}^{(n+(m-n-1))}\subsetneq F_{\text{general}}^{(n+(m-n-2))}\subsetneq\cdots\subsetneq F_{\text{general}}^{(n)}$
Appendix I Case Study
@@
I.1 Reaction Tasks
In this subsection, we visualize specific reactions in three reaction tasks. The results in Figure 12 and 13 demonstrate that our method exhibits more accurate generation capabilities compared to the baseline. For example, in the solvent prediction task, we are given the reactants: [C][C][Branch1][C][C][Branch1][C][C][O][C][=Branch1][C][=O][C][C][C][Br] and [C][C][=C][C][Branch1][Ring1][C][#N][=C][C][Branch1][C][Cl][=C][Ring1][=Branch2][O], as well as the product [C][C][=C][C][Branch1][Ring1][C][#N][=C][C][Branch1][C][Cl][=C][Ring1][=Branch2][O][C][C][C][C][=Branch1][C][=O][O][C][Branch1][C][C][Branch1][C][C][C]. Omni-Mol correctly predicts the solvent as [C][N][Branch1][C][C][C][=O], whereas PRESTO predicts an incorrect solvent: [C][C][#N].
I.2 Molecular Captioning
@@
Appendix J Task Definition & Prompt Templates
J.1 Base Chat Template
@@
J.2 Forward Reaction Prediction
The forward reaction prediction task focuses on determining the chemical product of a reaction given its reactants and reagents. The forward reaction prediction task involves predicting the chemical product of a reaction given the reactants and reagents as input. The input format is structured as the SELFIES representation of reactants, concatenated with a period (“.”) and the reagent information (e.g., “reactant1.reactant2.reagent”). The task requires the model to process this input and output the corresponding reaction product. The objective is to accurately map the input reaction components to their chemical outcome, leveraging the model’s understanding of reaction patterns and transformations. A key challenge in forward reaction prediction is capturing the underlying chemical rules that govern reactivity. The model must infer how functional groups interact, recognize the role of reagents, and apply appropriate transformations to generate the correct product. This process requires a deep understanding of reaction mechanisms, beyond simple pattern recognition. The prompt template is as follows.
@@
Instruction: Given the reactants and reagents provided, what is a possible product that can be formed? ¡SELFIES_reagents¿: [Br][C][C][Br].[O][C][=C][C][Branch1][C][Br][=C][C][=C][Ring1][#Branch1][Br] ¡SELFIES_reactants¿: [Na+1].[OH1-1] ¡SELFIES_product¿: [Br][C][C][O][C][=C][C][Branch1][C][Br][=C][C][=C][Ring1][#Branch1][Br]
J.3 Retrosynthesis
@@
Instruction: Which reactants could have been used to generate the given product? The product is: ¡SELFIES_product¿: [C][C][=Branch1][C][=O][C][=C][C][=C][Branch1][C][O][C][Branch1][C][Cl][=C][Ring1][Branch2] ¡SELFIES_reactants¿: [C][C][=Branch1][C][=O][Cl].[O][C][=C][C][=C][C][=C][Ring1][=Branch1][Cl]
J.4 Reagent Prediction
The reagent prediction task focuses on identifying the necessary reagents for a given chemical reaction, a critical step in reaction planning and synthetic chemistry. This task is essential for guiding experimental chemists, as choosing the correct reagents influences reaction efficiency, selectivity, and feasibility. To ensure a structured and standardized input format, we represent the reaction equation using SELFIES, a robust molecular encoding system. The input consists of the SELFIES representations of the reactants, concatenated with a reaction separator “>>”, followed by the SELFIES representation of the product. This format (e.g., “reactant1.reactant2>>product”) provides a clear, machine-readable structure that allows the model to infer the necessary reagents based on known reaction mechanisms and transformation rules. One of the core challenges in reagent prediction is handling the diversity of chemical transformations. Different reactions require specific reagents that dictate the reaction type, whether it’s an oxidation, reduction, coupling, or substitution reaction. The model must learn to recognize reaction context, interpret functional group interactions, and infer the most likely reagents from training data. The prompt template is as follows.
@@
Instruction: Can you provide potential reagents for the following chemical reaction? The reaction is ¡SELFIES_reactants¿: [C][C][C][Branch1][#C][C][=C][C][=N][C][Branch1][Ring1][O][C][=C][Ring1][Branch2][C][=O][O][C][C][C][O][Ring1][S] ¡SELFIES_product¿: [C][C][C][Branch1][#C][C][=C][C][=N][C][Branch1][Ring1][O][C][=C][Ring1][Branch2][C][O][O][C][C][C][O][Ring1][S] ¡SELFIES_reagents¿: [C][C][Branch1][C][C][O].[O].[BH4-1].[Na+1]
J.5 Molecular Captioning
The molecular captioning (Molcap) task focuses on generating descriptive textual information for a given chemical compound based on its molecular structure. This task plays a crucial role in chemical informatics, enabling automated annotation of molecular properties, classification, and functional characteristics. MolCap leverages machine learning models to infer and generate human-readable descriptions that encapsulate key chemical attributes. The input for this task follows a structured format using SELFIES, a robust molecular representation designed for machine learning applications. The SELFIES encoding of a given compound serves as the input, and the model is responsible for producing a descriptive caption that includes relevant chemical properties. These descriptions can encompass a wide range of molecular characteristics, such as compound classification (e.g., “organic acid,” “amine-containing molecule”), pH estimation, presence of functional groups (e.g., “contains a hydroxyl and ketone group”), solubility, toxicity, or other key features. One of the key challenges in molecular captioning is ensuring that the generated text is both chemically accurate and contextually informative. The model must learn to recognize molecular substructures, infer meaningful chemical attributes, and articulate these in a clear and interpretable manner. The prompt template is as follows.
@@
Instruction: Please give me some details about this molecule. The compound SELFIES sequence is: ¡SELFIES_compound¿:[C][=C][C][=Branch2][Ring1][S][=C][C][=C][Ring1][=Branch1][C][C][Branch2][Ring1][Branch1][C][Branch1][P][C] [Branch1][Ring2][O][Ring1][Branch1][C][O][P][=Branch1][C][=O][Branch1][C][O-1][O-1][O][O][O] Description: The molecule is an organophosphate oxoanion obtained by deprotonation of the phosphate OH groups of 4-(5-O-phospho-beta-D-ribofuranosyl)phenol; major species at pH 7.3. It derives from a D-ribofuranose 5-phosphate(2-). It is a conjugate base of a 4-(5-O-phospho-beta-D-ribofuranosyl)phenol.
J.6 Quantum Mechanics Property Prediction

### v3

@@ Appendix F More ablation study results
Model Exact BLEU Levenshtein RDK MACCS Morgan Validity
Forward Reaction Prediction Task
Head tuning 0.56 0.927 9.98 0.791 0.888 0.72 1.00
Omni-Mol 0.73 0.980 5.55 0.895 0.947 0.87 1.00
Model Exact BLEU Levenshtein RDK MACCS Morgan Validity
Retrosynthesis Task
Head tuning 0.45 0.942 12.82 0.79 0.865 0.75 1.00
Omni-Mol 0.57 0.960 8.97 0.864 0.909 0.83 1.00
Model Exact BLEU Levenshtein RDK MACCS Morgan Validity
Reagent Prediction Task
Head tuning 0.14 0.662 17.74 0.459 0.556 0.42 1.00
Omni-Mol 0.23 0.726 14.59 0.557 0.627 0.52 1.00
Model Exact BLEU Levenshtein RDK MACCS Morgan Validity
Catalyst Prediction
Head tuning 0.71 0.727 2.56 0.888 0.872 0.72 1.00
Omni-Mol 0.73 0.980 5.55 0.895 0.947 0.87 1.00
Model BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Molecular Captioning Task
Head tuning 0.450 0.338 0.546 0.359 0.477 0.482
Omni-Mol 0.529 0.440 0.604 0.447 0.541 0.571
Model Exact BLEU Levenshtein RDK MACCS Morgan Validity
Solvent Prediction
Head tuning 0.39 0.68 3.34 0.530 0.547 0.50 1.00
Omni-Mol 0.52 0.759 2.71 0.671 0.673 0.64 1.00
Table 9: Additional ablation results on 6 tasks. We freeze the LLM decoder and only activate the language model head.
F.1 Ablation on Language Backbone
To ensure that the performance gain doesn’t come from newer language backbones like Llama3.1 and Llama3.2, we replace the language model in Omni-Mol to Vicuna 7B [80] and conduct experience on several tasks. As shown in Table 11, with larger parameter size, Vicuna performs significantly better than Llama3.
F.2 Ablation on Parameter Efficient Tuning
One alternative tuning method is to tune only the language model head while keeping all parameters in the LLM decoder layers frozen. Note that for Llama 3.2 1B, the weights of the language model head are tied with the word embeddings; therefore, tuning the language model head also updates the word embeddings. The results are shown in Table 9, where head tuning performs significantly worse than Omni-Mol.
F.3 Ablation on Clip in GAL
Clipping is employed to prevent the model from converging to local optima during training. To validate its effectiveness, we remove the clipping mechanism in this experiment. As shown in Table 10, removing clip generally results in performance drop.
Method Exact BLEU RDK MACCS Morgan Levenshtein Validity
Catalyst Prediction
Omni-Mol 0.742 0.794 0.911 0.899 0.747 1.843 1.000
w/o clip 0.731 0.775 0.898 0.878 0.737 2.144 1.000
Forward Prediction
Omni-Mol 0.738 0.985 0.884 0.941 0.865 6.06 1.000
w/o clip 0.703 0.980 0.865 0.927 0.840 7.045 1.000
Reagent Prediction
Omni-Mol 0.266 0.749 0.586 0.651 0.542 14.026 1.000
w/o clip 0.256 0.736 0.577 0.645 0.533 14.619 1.000
Retrosynthesis
Omni-Mol 0.594 0.962 0.861 0.910 0.828 8.386 1.000
w/o clip 0.542 0.959 0.838 0.897 0.802 9.937 1.000
Solvent
Omni-Mol 0.554 0.782 0.703 0.700 0.678 2.498 1.000
w/o clip 0.550 0.775 0.692 0.693 0.667 2.550 1.000
Method HOMO LUMO GAP AVG
Quantum Mechanics
Omni-Mol 0.0038 0.0047 0.0049 0.0044
w/o clip 0.0039 0.0055 0.0047 0.0047
Method B-2 B-4 M R-1 R-2 R-L
Molcap
Omni-Mol 0.511 0.421 0.556 0.593 0.434 0.530
w/o clip 0.483 0.391 0.526 0.570 0.406 0.507
Method BH SM
Yield
Omni-Mol 0.953 0.688
w/o clip 0.933 0.680
Table 10: Ablation on clip in GAL
Model #Par Exa BLEU Lev RDK MAC Mor Val
Forward Reaction Prediction Task
Vicuna 7B [80] SL 6.7B 0.29 0.94 16.86 0.768 0.610 0.557 1.00
Llama3 1B [18] SL 1.3B 0.28 0.92 17.03 0.759 0.603 0.555 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
Solvent Prediction
Vicuna 7B [80] SL 6.7B 0.29 0.62 4.071 0.454 0.430 0.392 1.00
Llama3 1B [18] SL 1.3B 0.28 0.60 4.214 0.453 0.425 0.388 1.00
Model Type #Param BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Molecular Captioning Task
Vicuna 7B [80] SL 6.7B 0.483 0.390 0.559 0.386 0.493 0.522
Llama3 1B [18] SL 1.3B 0.455 0.361 0.549 0.381 0.487 0.503
Model #Par LogP
More Mol2Num
Vicuna 7B [80] SL 6.7B 0.5182
Llama3 1B [18] SL 1.3B 0.6690
Table 11: Comparison with different language backbone. We changed the backbone to Vicuna 7B and trained on several tasks, the results clearly indicate that better performance can be obtained from larger language model.
@@ F.3 Ablation on Clip in GAL
Appendix G Case Study
@@
Figure 9: Visualization of the cases generated by InstructMol and Omni-Mol on the molecule editing task.
G.1 Reaction Tasks
In this subsection, we visualize specific reactions in three reaction tasks. The results in Figure 7 and 8 demonstrate that our method exhibits more accurate generation capabilities compared to the baseline. For example, in the solvent prediction task, we are given the reactants: [C][C][Branch1][C][C][Branch1][C][C][O][C][=Branch1][C][=O][C][C][C][Br] and [C][C][=C][C][Branch1][Ring1][C][#N][=C][C][Branch1][C][Cl][=C][Ring1][=Branch2][O], as well as the product [C][C][=C][C][Branch1][Ring1][C][#N][=C][C][Branch1][C][Cl][=C][Ring1][=Branch2][O][C][C][C][C][=Branch1][C][=O][O][C][Branch1][C][C][Branch1][C][C][C]. Omni-Mol correctly predicts the solvent as [C][N][Branch1][C][C][C][=O], whereas PRESTO predicts an incorrect solvent: [C][C][#N].
G.2 Molecular Captioning
@@
G.3 Molecule Editing
The case study of the molecule editing task is presented in Figure 9. In this section, we analyze Omni-Mol’s “modus operandi” in molecule editing from a biological perspective, examining scaffold architecture, substituent patterns, and other insights to elucidate its potential real-world applicability.
We select optimized molecules meeting both unconstrained and constrained conditions and exceeding the QED threshold (0.6) for comparison, thereby revealing the optimization focus under each setting. And QED is modeled as the desirability profile of eight molecular properties, including molecular weight (MW: 250–400), lipophilicity (AlogP: 1–4), hydrogen bond donors (HBD: 0–2), hydrogen bond acceptors (HBA: 3–6), polar surface area (PSA: 40–90), number of rotatable bonds (ROTB < 10), aromatic rings (AROM: 0–2), and structural alerts (ALERTS: 0) [5].
For case 1, both InstructMol and Omni-Mol retain the benzenediazonium core scaffold while targeting the N-(3-methoxyphenyl)carbamimidothioic acid substituent, which features two hydrogen-bond donors, two acceptors, two rotatable bonds, and three structural alerts—most notably the carbamimidothioic acid (thiourea-like) moiety associated with acute toxicity. InstructMol merely replaces the terminal anisole with a 1,3-benzodioxole, failing to eliminate the toxicophore and producing negligible changes in overall properties (only one rotatable bond is removed), thus yielding a minimal QED improvement. By contrast, Omni-Mol replaces the entire N-(3-methoxyphenyl)carbamimidothioic acid substituent with 1-methylpiperazine, eliminating all alerts, removing an aromatic ring, substantially reducing molecular weight (from 479.6 to 397.5 Da) and rotatable bonds, and lowering H-bond donors to one. Benefiting from the hydrophilicity of the 1-methylpiperazine substituent, ALOGP is markedly decreased into the developable range.
For case 2, under unconstrained conditions (no hard similarity requirement), the optimization is more flexible, better revealing Omni-Mol’s design logic versus InstructMol. InstructMol’s molecule still contains multiple aromatic rings, whereas Omni-Mol removes excess aromatics, preserves only the core scaffold, and introduces small saturated rings (e.g., cyclopropyl) to enhance three-dimensionality and rigidity, mitigating risks of over-aromatization. InstructMol’s substituents include heavy‐atom hydrophobic groups such as trifluoromethyl and fluoro on the aromatic ring, and a reactive aldehyde (CHO-CHO) at the terminus. Omni-Mol, however, introduces an ethoxy group in place of the $CF_{3}$ unit and installs an amide–cyclopropane fragment, improving physicochemical balance and stability. For side-chain optimization, instead of InstructMol’s long, highly polar chain, Omni-Mol cleverly replaces the bulky alkyl-alcohol side chain with a small, conformationally constrained cyclopropyl substituent, also boosting lipophilicity. These combined effects raise the QED score.
In summary, Omni-Mol exhibits a flexible yet systematic optimization strategy—scaffold tuning, side-chain simplification, and substituent “play to strengths and avoid weaknesses”, achieving moderate molecular size and balanced hydrophobicity. In contrast, InstructMol tends toward rigid expansion and structural complexity, lacking an adaptable, cohesive design rationale. Thus, Omni-Mol’s QED gains derive from rational, streamlined principles that balance flexibility and systemacity, aligning with medicinal chemistry guidelines and reflecting a design ethos likely to resonate with real-world biochemical researchers, offering substantial practical value and significance. More detailed information about the QED properties can be found in Table 12.
Model MW ALOGP HBA HBD PSA ROTB AROM ALERTS
QED Properties for Case 1
Original 479.60 6.08 5 3 81.18 9 3 3
InstructMol 493.59 5.80 6 3 90.41 8 3 3
Omni-Mol 397.52 3.80 5 1 54.04 7 2 0
QED Properties for Case 2
Original 605.59 6.06 6 0 90.21 9 5 1
InstructMol 587.62 4.85 7 2 98.66 11 3 2
Omni-Mol 434.54 4.27 5 0 67.26 8 1 0
Table 12: Comparison of QED Properties between InstructMol, and Omni-Mol for two cases
Appendix H task definition and prompt templates
H.1 Base Chat Template
@@
H.2 Forward Reaction Prediction
The forward reaction prediction task focuses on determining the chemical product of a reaction given its reactants and reagents. The forward reaction prediction task involves predicting the chemical product of a reaction given the reactants and reagents as input. The input format is structured as the SELFIES representation of reactants, concatenated with a period (“.") and the reagent information (e.g., “reactant1.reactant2.reagent"). The task requires the model to process this input and output the corresponding reaction product. The objective is to accurately map the input reaction components to their chemical outcome, leveraging the model’s understanding of reaction patterns and transformations. A key challenge in forward reaction prediction is capturing the underlying chemical rules that govern reactivity. The model must infer how functional groups interact, recognize the role of reagents, and apply appropriate transformations to generate the correct product. This process requires a deep understanding of reaction mechanisms, beyond simple pattern recognition. The prompt template is as follows.
@@ H.2 Forward Reaction Prediction
Instruction: Given the reactants and reagents provided, what is a possible product that can be formed? <SELFIES_reagents>: [Br][C][C][Br].[O][C][=C][C][Branch1][C][Br][=C][C][=C][Ring1][#Branch1][Br] <SELFIES_reactants>: [Na+1].[OH1-1] <SELFIES_product>: [Br][C][C][O][C][=C][C][Branch1][C][Br][=C][C][=C][Ring1][#Branch1][Br]
H.3 Retrosynthesis
@@
Instruction: Which reactants could have been used to generate the given product? The product is: <SELFIES_product>: [C][C][=Branch1][C][=O][C][=C][C][=C][Branch1][C][O][C][Branch1][C][Cl][=C][Ring1][Branch2] <SELFIES_reactants>: [C][C][=Branch1][C][=O][Cl].[O][C][=C][C][=C][C][=C][Ring1][=Branch1][Cl]
H.4 Reagent Prediction
The reagent prediction task focuses on identifying the necessary reagents for a given chemical reaction, a critical step in reaction planning and synthetic chemistry. This task is essential for guiding experimental chemists, as choosing the correct reagents influences reaction efficiency, selectivity, and feasibility. To ensure a structured and standardized input format, we represent the reaction equation using SELFIES, a robust molecular encoding system. The input consists of the SELFIES representations of the reactants, concatenated with a reaction separator “>>", followed by the SELFIES representation of the product. This format (e.g., “reactant1.reactant2>>product") provides a clear, machine-readable structure that allows the model to infer the necessary reagents based on known reaction mechanisms and transformation rules. One of the core challenges in reagent prediction is handling the diversity of chemical transformations. Different reactions require specific reagents that dictate the reaction type, whether it’s an oxidation, reduction, coupling, or substitution reaction. The model must learn to recognize reaction context, interpret functional group interactions, and infer the most likely reagents from training data. The prompt template is as follows.
@@ H.4 Reagent Prediction
Instruction: Can you provide potential reagents for the following chemical reaction? The reaction is <SELFIES_reactants>: [C][C][C][Branch1][#C][C][=C][C][=N][C][Branch1][Ring1][O][C][=C][Ring1][Branch2][C][=O][O][C][C][C][O][Ring1][S] <SELFIES_product>: [C][C][C][Branch1][#C][C][=C][C][=N][C][Branch1][Ring1][O][C][=C][Ring1][Branch2][C][O][O][C][C][C][O][Ring1][S] <SELFIES_reagents>: [C][C][Branch1][C][C][O].[O].[BH4-1].[Na+1]
H.5 Molecular Captioning
The molecular captioning (Molcap) task focuses on generating descriptive textual information for a given chemical compound based on its molecular structure. This task plays a crucial role in chemical informatics, enabling automated annotation of molecular properties, classification, and functional characteristics. MolCap leverages machine learning models to infer and generate human-readable descriptions that encapsulate key chemical attributes. The input for this task follows a structured format using SELFIES, a robust molecular representation designed for machine learning applications. The SELFIES encoding of a given compound serves as the input, and the model is responsible for producing a descriptive caption that includes relevant chemical properties. These descriptions can encompass a wide range of molecular characteristics, such as compound classification (e.g., “organic acid," “amine-containing molecule"), pH estimation, presence of functional groups (e.g., “contains a hydroxyl and ketone group"), solubility, toxicity, or other key features. One of the key challenges in molecular captioning is ensuring that the generated text is both chemically accurate and contextually informative. The model must learn to recognize molecular substructures, infer meaningful chemical attributes, and articulate these in a clear and interpretable manner. The prompt template is as follows.
@@ H.5 Molecular Captioning
Instruction: Please give me some details about this molecule. The compound SELFIES sequence is: <SELFIES_compound>:[C][=C][C][=Branch2][Ring1][S][=C][C][=C][Ring1][=Branch1][C][C][Branch2][Ring1][Branch1][C][Branch1][P][C] [Branch1][Ring2][O][Ring1][Branch1][C][O][P][=Branch1][C][=O][Branch1][C][O-1][O-1][O]$[O]$ Description: The molecule is an organophosphate oxoanion obtained by deprotonation of the phosphate OH groups of 4-(5-O-phospho-beta-D-ribofuranosyl)phenol; major species at pH 7.3. It derives from a D-ribofuranose 5-phosphate(2-). It is a conjugate base of a 4-(5-O-phospho-beta-D-ribofuranosyl)phenol.
H.6 Quantum Mechanics Property Prediction